### v2

# Alexandra Ciobotaru

Abstract
Psychological corpora in NLP are collections of texts used to analyze human psychology, emotions, and mental health. These texts allow researchers to study psychological constructs, detect mental health issues and analyze emotional language. However, mental health data can be difficult to collect correctly from social media, due to suppositions made by the collectors. A more pragmatic strategy involves gathering data through open-ended questions and then assessing this information with self-report screening surveys. This method was employed successfully for English, a language with a lot of psychological NLP resources. However, this cannot be stated for Romanian, which currently has no open-source mental health corpus. To address this gap, we have created the first corpus for depression and anxiety in Romanian, by utilizing a form with 6 open-ended questions along with the standardized PHQ-9 and GAD-7 screening questionnaires. Consisting of the texts of 205 respondents and although it may seem small, PsihoRo is a first step towards understanding and analyzing texts regarding the mental health of the Romanian population. We employ statistical analysis, text analysis using Romanian LIWC, emotion detection and topic modeling to show what are the most important features of this newly introduced resource to the NLP community.
Keywords mental health $\cdot$ Romanian $\cdot$ emotion detection $\cdot$ LIWC $\cdot$ topic modeling
1 Introduction
Mental health research has increasingly used data from social media platforms, which often present a curated view of individuals’ personal lives. While social media can provide valuable insights for mental health studies [22, 27], there is a risk of positivity bias, where users tend to share mostly positive aspects of their experiences [29]. Another important source for mental health research is clinical data, which typically contains sensitive information and is not publicly accessible.
Most research on mental health is conducted using data in English, which may overlook important nuances present in other languages. Previous studies suggest that features predictive of mental health issues in English may not translate effectively to other languages or cultures [7, 27]. For example, the use of the first-person pronoun "I" has been identified as a strong predictor of depression [28], and this feature is frequently analyzed in English-based research. However, the relationship between pronoun usage and depression severity can vary significantly across different demographic groups and languages [27]. While the pronoun "I" is often used as an indicator of depression in English, its usage requires special consideration in other languages due to linguistic differences. For instance, previous research has shown that this association has not been observed among Romanian speakers [33]. Romanian is a pro-drop language, which allows for ambiguity in subject pronoun usage and may contribute to a lower frequency of the personal pronoun ”I” in everyday communication.
Despite the availability of some data in the Romanian language related to emotions, such as the RED dataset [9], REDv2 [8], and the BRIGHTER multilingual dataset [24], which includes Romanian data, there is a scarcity of mental health-related resources. Very few clinical data sources [33] and posts from medical forums [6, 11] exist, and those that do are not publicly accessible.
This study aims to address this gap by creating the first publicly available dataset in Romanian, specifically designed to analyze emotional expressions in the context of depression and anxiety. By analyzing the linguistic characteristics and emotional patterns in Romanian, our research will contribute to a deeper understanding of mental health expressions in languages other than English. Through this initiative, we seek to promote more inclusive research in mental health, paving the way for targeted interventions and support across diverse linguistic communities.
2 Related Work
[35] created open-ended personality questions to explore how natural language processing can aid in predicting neuroticism from people’s responses. They gathered over 30,000 sentences written by 425 adults in Korean and used KoBERT to see which types of questions best revealed neuroticism and related traits like depression and dependency.
[34] created the multi-modal Real World Worry Waves Dataset (RW3D), in English, which tracks the psychological effects of the COVID-19 pandemic in the UK over three consecutive years, 2020, 2021 and 2022, through repeated surveys containing also free-text responses.
[10] used open-ended interview responses from more than 58,000 job candidates along with self-assessed personality (HEXACO) and employed “InterviewBERT” to deduce personality traits from textual replies. This study was also done in English.
[30] used open-ended questionnaires to study young and middle aged adults’ responses regarding their own mental health, by employing NLP computational methods. They analyzed their 1106 responses and found various differences in semantics on the two study groups.
[3] combined traditional linguistic analysis (LIWC and sentiment analysis) with NLP and machine learning to find a relationship between text style and depression, anxiety, and stress during the COVID-19 pandemic. They analyzed written narratives from 883 participants combined with their DASS-21 evaluation scores and found a connection between increased symptom severity and the greater use of first-person pronouns and negative emotion words.
[19] collected responses in English from 411 participants who answered word-response questions related to their mental health and completed also the GAD-7 and PHQ-9 rating scales. Their analyses conclude that individuals’ language reflects their mental health status, and combining verbal responses with quantitative measures can enhance the comprehension and assessment the risk of depressive and anxiety symptoms in both clinical practice and research contexts.
In terms of combing emotion detection and LIWC term analysis to mental health texts, there are several studies which employ such analysis on the English language: [26], [36], [2].
There are several datasets created by combining gathering text through open ended questionnaires together with a screening tool for detecting the risk of a potential mental problem, but mainly centered on the English language. For Romanian, currently there are no open source psychological corpora available, and this fact was the main driving force behind our study, along with the curiosity of applying emotional analysis on such a corpus.
3 Data Collection
By using this survey collection methodology, participants provide their informed consent for the collection of data, as opposed to social media collected data, where data is collected without asking the participants for their consent [4].
The data collection methodology consists of a completely anonymous survey, organized in three parts. The first part of the survey consists of six open-ended questions. Its scope is to accumulate as much written text as possible. Out of these six questions, three of them touch positive topics and three of them negative topics. The survey was written in Romanian for Romanian speakers only. Below are English translations for the six open-ended questions:
  1. 1.
    To what extent do you think social media affects your emotional state, especially when you see negative news? "În ce măsură crezi că rețelele sociale îți afectează starea emoțională, mai ales când vezi știri negative?" (negative)
  2. 2.
    What is an important lesson about resilience and adaptation that you have learned recently? "Care este o lecție importantă despre reziliență și adaptare pe care ai învățat-o recent?" (positive)
  3. 3.
    How much do you think politics affects people’s mental health? Can you describe a time when you felt stressed because of political decisions? "Cât de mult crezi că politica afectează sănătatea mentală a oamenilor? Poți să detaliezi un moment în care ai simțit stres din cauza deciziilor politice?" (negative)
  4. 4.
    What role do you think community or social support plays in maintaining mental health during difficult times? "Ce rol crezi că are comunitatea sau sprijinul social în menținerea sănătății mentale în vremuri dificile? "(positive)
  5. 5.
    How were you affected by discussions about war, the economic crisis, elections, or other tense topics? "Cum ai fost afectat de discuțiile despre război, criza economică, alegeri sau alte subiecte tensionate?" (negative)
  6. 6.
    How do you recharge your batteries after a stressful period? "Cum îți reîncarci bateriile după o perioadă stresantă?" (positive)
We wanted to create a corpus of text data written by persons prone to suffering from depression and/or anxiety. Usually psychological corpuses for NLP are usually gathered from social media through searching searching different keywords in social media posts that could indicate depression and/or anxiety. By discussing with medical professionals in the psychiatric field we concluded that this type of gathering psychological data is speculative. Thus, we added to our survey two more parts, one that assesses if the respondent is prone to depression by incorporating PHQ-9 [21] questions, and another part which assesses if the respondent is prone to anxiety by incorporating GAD-7 [32] questions. Both GAD-7 and PHQ-9 contain questions (7 for GAD-7 and 9 for PHQ-9) which ask how often the respondent has been bothered by certain problems, and the response options correspond to a numerical score: Not at all = 0, Several days = 1, More than half the days = 2, Nearly every day = 3.
The severity cut-offs for depression based on PHQ-9 are: 0-4: Minimal, 5-9: Mild, 10-14: Moderate, 15-19: Moderately severe, 20-27: Severe, while the severity cut-offs for anxiety based on GAD-7 are: 0-4: Minimal, 5-9: Mild, 10-14: Moderate, 15-19: Moderately severe, 20-27: Severe.
The resulting questionnaire is completely anonymous, no personal data is requested or stored (such as name, surname, email address, IP). We did not collect demographic data for the same reason: to gain the trust of respondents and also to keep the total time of form completion short.
The resulting data consists of 205 responses, collected over an eight-month period from March 2025 to October 2025. This time frame includes two major stressful events for Romanians, presidential elections and the tax increase burden generated by the country’s new leadership.
4 Methodology
Our research consists of gathering insights from PsihoRo descriptive statistics in correlation to depression and anxiety respondents’ tendencies, for three main population groups which were computed based on the severity cut-offs. According to the literature in psychology [20], a cut-off score of 10 or higher on the PHQ-9 questionnaire is used to identify the presence of potential major depressive disorder, prompting further assessment. For the GAD-7 questionnaire, also a cut-off score of 10 has been proposed as a threshold for identifying potential generalized anxiety disorder, warranting further evaluation. [31].
Based on these rules, three population groups were derived from our dataset:
  • Depression Risk Group: 37 respondents (18%) scored higher than 10 on PHQ-9;
  • Anxiety Risk Group: 36 respondents (17.56%) scored higher than 10 on GAD-7;
  • No Risk Group: 28 respondents (13.66%) scored 0 for at least one out of the two questionnaires.
Linguistic Inquiry and Word Count (LIWC) [5] is a lexicon and a text analysis software that provides psychologically meaningful categories and is widely used in computational mental health research [28, 22, 15, 33]. The official LIWC software is available for analyzing textual data in the English language; however, validated translations of the software are also available in several languages.
In our experiments, we use the Romanian version of LIWC (Ro-LIWC2015) [11]. The 2015 version of LIWC includes 98 categories that relate to linguistic dimensions and grammar, as well as psychological processes, including affective, cognitive, social, perceptual, and biological processes.
Previous research has indicated that individuals with mental health issues tend to exhibit specific language patterns. These patterns include a higher frequency of using the personal pronoun "I" [28], increased use of words with negative connotations [13], and a tendency to use words with absolute meanings [1]. These findings highlight the importance of language analysis in understanding mental health and motivate us to use LIWC to examine the language patterns of the participants. After checking for differences in LIWC categories in relation to the PHQ-9 and GAD-7 scores, we wanted to test their discriminatory power in classification. To this end, we employed a LightGBM classifier [18] and trained it on the results with LIWC to predict a binary label for depression and anxiety, based on the cutoff score of 10. The model achieved an accuracy of 0.85 in classifying depression and 0.83 in classifying anxiety. To determine which LIWC features have the greatest influence on predictions, we used SHapley Additive exPlanations (SHAP) [23].
In parallel to these anlyses we have created an emotion detection model by fine-tuning the Romanian BERT [12] on the REDv2 dataset [8], for multi-label text classification, using the Huggingface transformers pipeline [^1] https://huggingface.co/docs/transformers. We used this model to predict the emotions in PsihoRo, thus gathering emotional information for the text written by each population group mentioned above. The resulting model had an F1 score of 66.85% computed on the REDv2 test set.
Another interesting analysis was topic modeling. To identify the main themes in the PsihoRo corpus we used the stm R package. First, the texts were tokenized, lowercased, and cleaned by removing punctuation, numbers, symbols and stopwords. Next, we created a Document-Feature Matrix (DFM) to represent term frequencies across texts. Also, to enhance model performance and reduce noise, we excluded rare terms which appeared in fewer than 3 documents and the very common terms, which appeared in more than 90% of the texts. Further, the cleaned DFM was transformed into the format necessary for the Structural Topic Model (STM) using which we trained the model for K = 5 topics, with the Spectral initialization method.
In order to gain interpretable results, the words resulted from the most important topics were further expanded into their main reformulated ideas using GPT-5.
5 Experiments and Results
5.1 PsihoRo Descriptive Statistics
In Table 1 we show mean, median (50% quantile), min, max, 25% and 75% cuantiles, and the standard deviation for both anxiety and depression scores. It can be seen that the maximum value for GAD-7 (anxiety) is 21, which is the maximum which can be achieved, and the maximum score for PHQ-9 (depression) is 23, out of the maximum possible of 27.
Table 1: Descriptive statistics for PsihoRo
PHQ-9 GAD-7
count 205 205
mean 6 5.78
standard deviation 4.83 4.95
min 0 0
25% 3 2
50% (median) 5 5
75% 8 7
max 23 21
By analyzing the length of the answers, we questioned whether people with risk of depression and/or anxiety tend to write less. The answer to this question is no, because the computed T-statistic was 1.55 between the depression risk population and the no risk population, and 2.26 between the anxiety risk population and the no risk population, meaning that no significant difference was found.
In Figure 1 we show the distribution plots of both PHQ-9 and GAD-7 scores. The data showed that the distributions are skewed to the left, with the majority of observations concentrated at lower values and a longer tail extending toward the higher end of the scale, meaning that most respondents have minimal and mild risk of depression and anxiety.
Figure 1: Distributions for GAD-7 and PHQ-9
The Pearson correlation coefficient (r) between PHQ-9 and GAD-7 total score is 0.761 (p<0.001), which means that anxiety and depression are strongly correlated in our data. This result shows as well in Figure 2 and is in concordance with the results of several other works: [17], [14], [16].
Figure 2: Correlation between Anxiety and Depression Scores.
5.2 LIWC
We employ LIWC to analyze language patterns in participants’ open-ended responses and correlate them with scores from the PHQ-9 and GAD-7. In Table 2, we present the Pearson correlation coefficients between LIWC categories and the symptom scores. Positive values indicate that higher symptom scores are associated with greater use of that linguistic category, while negative values indicate the opposite. Asterisks indicate significance (*p < .05, **p < .01).
Surprisingly, the use of negations is negatively correlated with both PHQ-9 and GAD-7 scores, suggesting that participants with higher depression and anxiety scores tend to use fewer negations in their responses. Individuals who score higher on these assessments tend to use more words in the Tentative category, which indicates uncertainty, supported by the correlation coefficients of .204** for PHQ-9 and .213** for GAD-7. Comparatives also show a positive correlation with the depression score (.167*), indicating comparative thinking, which is often linked to depressive rumination. However, another marker for rumination, the high frequency of Past focus verbs, is not observed in our data. There is a positive correlation between the use of Body-related words and PHQ-9 (200**) and GAD-7 (180**) scores, indicating that participants with higher scores frequently use more language related to bodily sensations. This may reflect increased somatic awareness, a common trait in depression and anxiety [25]. Participants with higher PHQ-9 and GAD-7 scores discuss Achievement less frequently (-.163* and- .192**, respectively), which could indicate low motivation. In addition, the negative correlation with Leisure-related words (-.145* and -.169*, respectively) may stem from anhedonia, a common symptom of depression. Words in the Death category show a significant positive correlation with depression (.138*), but not with anxiety (0.029), indicating a greater preoccupation with mortality in participants with higher depression scores. In contrast, the Family category is positively correlated with anxiety scores (r = 0.145*), suggesting that individuals with higher anxiety are more focused on family-related issues.
Table 2: Correlations between PHQ-9, GAD-7, and LIWC dimensions
LIWC Dimension Examples PHQ-9 GAD-7
I I, me, mine 0.018 -0.046
We we, us, our 0.013 -0.039
You you, your 0.006 0.084
Verb eat, come, carry -0.087 -0.091
Past focus ago, did, talked -0.064 -0.075
Present focus today, is, now -0.019 -0.034
Future focus may, will, soon -0.027 -0.055
Adjectives free, happy, long 0.118 0.12
Negations no, not, never -.156* -.167*
Positive emotion love, nice, sweet -0.122 -0.057
Negative emotion hurt, ugly, nasty 0.028 0.028
Anxiety worried, fearful -0.042 0.01
Anger hate, kill, annoyed 0.107 0.105
Sadness crying, grief, sad 0.074 0.045
Family daughter, dad, aunt 0.079 .145*
Friends buddy, neighbor -0.026 0.043
Female references girl, her, mom 0.061 0.074
Male references boy, his, dad .156* .177*
Comparatives greater, best, after .167* 0.081
Causation because, effect -.186** -.180**
Tentative maybe, perhaps .204** .213**
Insight think, know -0.075 -0.101
Assent agree, OK, yes 0.009 0.065
Perceptual processes look, heard, feeling 0.009 0.033
Body cheek, hands, spit .200** .180**
Health clinic, flu, pill -0.051 0.035
Achievement win, success, better -.163* -.192**
Leisure cook, chat, movie -.145* -.169*
Death bury, coffin, kill .138* 0.029
Note. * $p<.05$, ** $p<.01$.
Figure 3: Differences in selected LIWC linguistic categories between participants with low vs. high PHQ-9 (left) and GAD-7 (right) scores. * denotes statistically significant differences ($p<0.05$) measured using the Mann-Whitney U test. Positive values indicate greater relative use among participants with higher symptom scores.
To better understand the differences between higher and lower symptom scores for depression and anxiety, we conducted an additional analysis of the LIWC categories. According to the cutoff values for both the PHQ-9 and GAD-7 assessment tools, we define a low PHQ score as anything less than 10, while a score greater than 10 is classified as a high PHQ-9 score. In Figure 3, the goal was to identify linguistic markers that differentiate individuals with higher vs. lower symptom scores. Statistically significant differences are indicated with an asterisk (*), as determined by the Mann-Whitney U test. Positive values indicate greater relative use among participants with higher symptom scores, while negative values indicate greater use among low-symptom participants. The analysis produced results consistent with the correlation analysis, particularly regarding Achievement and tentativeness in relation to depression and anxiety. However, for depression, we observed a decreased use of Positive emotion words and an increased usage of Sadness words, which aligns with existing literature in psychology [33]. In addition, the language of participants with high GAD-7 scores showed a higher frequency of words in the Interrogatives category, indicating a tendency to question or seek clarity.
Figure 4: Feature importance for LIWC categories analyzed using SHAP.
To determine which LIWC features have the greatest influence on predictions, we used SHapley Additive exPlanations (SHAP) [23], as explained in Section 4, with the results presented in Figure 4. The importance of the top 20 features aligns with our correlation analysis and highlights the differences between low and high scores for both depression and anxiety.
The LIWC analyses presented in this section demonstrate that lexical indicators can effectively capture underlying psychological patterns related to depression and anxiety. However, the most important feature used in mental health research, the pronoun ”I”, appears to be less useful for the Romanian language, likely because it is a pro-drop language. This finding is consistent with the work of [33], which also indicated that this marker does not generalize well in a sample of patients with major depression.
5.3 Emotion Analysis
Using the emotion detection model described in Section 4, we predicted the emotions in all the texts of PsihoRo, and then we aggregated the results by using radar charts to visualize the overall emotional patterns. In Figure 5 we show the emotion predictions across all PsihoRo texts, and also for the pessimistic questions (1, 3, 5) and the optimistic ones (2, 4, 6).
Figure 5: Emotion predictions across: all PsihoRo texts (left), pessimistic questions (middle), and optimistic questions (right).
Surprise is an emotion that is completely absent from the PsihoRo data. The following emotions are more common, in sequence of appearance: anger, joy, fear, sadness, trust, and the most common, neutral. The emotional pattern for the more pessimistic questions is displayed in the center, while the emotional pattern for the more optimistic questions is displayed on the right. Emotions such as sadness and anger are on the rise on the left, while joy and trust are more prevalent on the right.
The emotions present in the texts written by the three primary demographic groups mentioned at the start of Section 4 were also examined in Figure 6. It can be seen that besides the neutral, which is strongly present across all groups
Figure 6: Emotion predictions across the three studied population groups.
We also tried to create several regression models, based on the textual PsihoRo data gathered and the GAD-7 and PHQ-9 scores, using the Romanian BERT for text vectorization and also adding emotional information as features, but all our experiments had negative R2.
5.4 Topic Modeling
We applied the topic modeling analysis described in Section 4 for each of the three control groups studied in this article. The main topics are derived from the six questions asked the form. However, some differences appear between the three groups of respondents, which can be found in the following subsections [^2] The following subsections were generated using GPT-5, based on the words generated by the topic modeling operation, some of which present in text in bold..
5.4.1 No Risk Group
Topic 1. This theme highlights how individuals perceive the current political and social period in a critical and sometimes negative way. The words “negative”, “decisions”, “politics”, “measure”, “news” and “period” indicate a concern for the effects of government decisions and how they are reflected in the media. The presence of the term “sport” suggests a possible comparison between different areas of public life.
Topic 2. The theme reflects the complex interaction between information conveyed through news, political life and how they influence the general state of people. Terms such as “politics”, “news”, “state”, “people” and “politics” indicate an increased attention to how media discourse shapes collective perceptions and emotions.
Topic 3. This topic addresses how individuals interpret and process information related to elections and political decisions. The terms “measure”, “depends”, “people”, “time”, “elections” and “example” suggest a personal analysis of the electoral context and an attempt to understand the relationship between information, time and decision.
Topic 4. The theme highlights the importance of authentic human interaction and the balance between the inner and social dimensions of the individual. The words "listen", "outer", "inner", "community" and "matter" suggest an openness to empathetic communication and a valorization of social belonging.
Topic 5. This topic focuses on the emotional dimension of everyday experiences, indicating how individuals manage their affective states and reactions in common situations. Words such as “affect”, “try”, “day”, “go” and “enough” express a constant effort to adapt and maintain psychological balance in the face of daily pressures by self regulation and personal reflection.
5.4.2 Depression Risk Group
Topic 1. This theme captures the perception of fatigue, daily routine and the constant effort that individuals put into their personal and professional lives. Associated terms such as “sleep”, “hard”, “walking” and “people” indicate a general state of exhaustion and the need for balance between personal and social life.
Topic 2. This theme highlights the importance of psychological resilience and the ability to adapt to stressful or unpredictable situations. Words such as “resilience”, “situation”, “time”, “community” and “important” suggest a solution-oriented approach and maintaining a state of balance in contexts of uncertainty.
Topic 3. The topic focuses on how individuals relate to political processes and the electoral climate. The terms “politics”, “elections”, “people”, “moment” and “listening” indicate both interest in political events and their influence on social attitudes and conversations.
Topic 4. This topic captures the potentially harmful impact of social networks on interpersonal relationships and community cohesion. The presence of the words “negative”, “social”, “networks”, “community” and “socialization” underlines the perception of a degradation of authentic communication and well-being caused by the digital environment.
Topic 5. This theme reflects the concern for affective states and how social and economic factors contribute to the intensification of collective anxiety. The terms “feel”, “affect”, “cause”, “anxiety”, “country” and “family” suggest an emotional perspective on everyday experiences and the general social context.
5.4.3 Anxiety Risk Group
Topic 1. This topic captures how major events, such as war, economic crises or socio-political changes, affect the perception of daily life. Words such as “affect”, “crisis”, “war”, “elections”, “life” and “time” indicate the concern for external instability and its influence on personal balance.
Topic 2. This topic highlights the close connection between political factors and individual emotional state. Terms such as “politics”, “stress”, “affect”, “feel”, “cause” and “family” suggest a direct relationship between socio-political events and the level of anxiety or psychological tension felt.
Topic 3. This topic addresses the role of community and collective resilience in the face of the constant flow of negative information from the media. The words “resilience”, “news”, “negative”, “community”, “important” and “trying” indicate an attempt to find the balance between information and protecting one’s mental state.
Topic 4. This topic explores the effects of social media on psychological state, with a focus on anxiety and social pressure. Terms such as “social”, “anxiety”, “state”, “hard”, “networks” and “present” suggest a connection between the intensive use of social networks and the degradation of emotional state. The topic reflects how virtual interactions can accentuate psychological discomfort and feelings of isolation.
Topic 5. This theme emphasizes the connection between people, community and the sense of belonging. The words “people”, “community”, “listen”, “dear” and “happen” highlight the importance of close relationships and solidarity in maintaining emotional balance.
6 Conclusion and Future Work
PsihoRo is the first open-source dataset for mental health in Romanian. It was created by utilizing a compound form containing both open ended questions for text gathering and the PHQ-9 and GAD-7 consecrated questionnaires to asses the risk of depression and/or anxiety of the respondents. Even if the dataset has only 205 entries, some very interesting insights can be observed.
Regarding the psychologically meaningful categories from LIWC, we observed differences in categories related to body-related words, which may indicate somatic awareness, and tentative words, which indicate uncertainty, that correlated positively with both depression and anxiety scores. A higher preoccupation with mortality was also observed in relation to depression, while anxiety scores were positively correlated with a preoccupation with family. While LIWC analyses showed that lexical indicators can effectively capture underlying psychological patterns related to depression and anxiety, we did not find any association between the linguistic marker ”I” and depression or anxiety in Romanian language, consistent with previous research.
The emotion analysis showed the prevalence of sadness emotions in the population group with risk of depression, as well as fear for the risk of anxiety population group, which is also in concordance with previous research.
The inability to develop a regression model with consistent results based on PsihoRo can be attributed both to the low number of samples and also to the naivety of assuming that depression and anxiety can be reliably discerned from texts, even when supplemented by self-assessment ratings like PHQ-9 and GAD-7. Depression and anxiety require meticulous analysis from multiple perspectives, rather than solely relying on an individual’s textual expression. However, by offering high-quality data with labels for anxiety and depression, PsihoRo seeks to support this intricate analysis.
7 Ethical Considerations and Limitations
The data collection was approved by The Ethics Committee of the University of Bucharest. All respondents provided their informed consent before answering the survey, which was completely anonymous, no personal data was requested or stored (such as name, surname, email address, IP, etc.).
Adding more data by increasing the number of respondents, by using crowdfunding or similar platforms, could greatly improve PsihoRo and we aim to do this in future versions of this resource.
Acknowledgments
The authors would like to thank psychologist Virginia Neacșu for her valuable feedback and suggestions throughout the development of this work.
This research was supported by the University of Bucharest and Druid AI.

### v3

@@
 However, this cannot be stated for Romanian, which currently has no open-source mental health corpus.
-To address this gap, we have created the first corpus for depression and anxiety in Romanian, by utilizing a form with 6 open-ended questions along with the standardized PHQ-9 and GAD-7 screening questionnaires.
+To address this gap, we have created the first open-source corpus for depression and anxiety in Romanian, by utilizing a form with 6 open-ended questions along with the standardized PHQ-9 and GAD-7 screening questionnaires.
 Consisting of the texts of 205 respondents and although it may seem small, PsihoRo is a first step towards understanding and analyzing texts regarding the mental health of the Romanian population.
 We employ statistical analysis, text analysis using Romanian LIWC, emotion detection and topic modeling to show what are the most important features of this newly introduced resource to the NLP community.
+The data is publicly available at https://huggingface.co/datasets/Alegzandra/PsihoRo.
 Keywords mental health $\cdot$ Romanian $\cdot$ emotion detection $\cdot$ LIWC $\cdot$ topic modeling
@@
 2 Related Work
+Previous research has gathered open-ended responses from participants to explore how expressions of mental health influence language.
 [39] created open-ended personality questions to explore how natural language processing can aid in predicting neuroticism from people’s responses.
@@
 They analyzed their 1106 responses and found various differences in semantics on the two study groups.
-[3] combined traditional linguistic analysis (LIWC and sentiment analysis) with NLP and machine learning to find a relationship between text style and depression, anxiety, and stress during the COVID-19 pandemic.
-They analyzed written narratives from 883 participants combined with their DASS-21 evaluation scores and found a connection between increased symptom severity and the greater use of first-person pronouns and negative emotion words.
-[19] collected responses in English from 411 participants who answered word-response questions related to their mental health and completed also the GAD-7 and PHQ-9 rating scales.
-Their analyses conclude that individuals’ language reflects their mental health status, and combining verbal responses with quantitative measures can enhance the comprehension and assessment the risk of depressive and anxiety symptoms in both clinical practice and research contexts.
-In terms of combing emotion detection and LIWC term analysis to mental health texts, there are several studies which employ such analysis on the English language: [26], [36], [2].
-There are several datasets created by combining gathering text through open ended questionnaires together with a screening tool for detecting the risk of a potential mental problem, but mainly centered on the English language.
-For Romanian, currently there are no open source psychological corpora available, and this fact was the main driving force behind our study, along with the curiosity of applying emotional analysis on such a corpus.
+In terms of combining emotion detection and LIWC term analysis in mental health texts, several studies employ such analysis in English: [29], [40], [2].
+[4] combined LIWC features [6], sentiment analysis and NLP to explore the relationship between language and mental health issues such as depression, anxiety, and stress during the COVID-19 pandemic.
+They analyzed written narratives from 883 participants, alongside their DASS-21 scores, and found an association between increased symptom severity and greater use of first-person pronouns and negative emotion words.
+[21] collected responses in English from 411 participants who answered word-response questions related to their mental health and completed the GAD-7 and PHQ-9 rating scales.
+Their analyses show that individuals’ language reflects their mental health status, and combining verbal responses with quantitative measures can enhance the comprehension and assessment of the risk of depressive and anxiety symptoms in both clinical practice and research contexts.
+While several datasets have been created using text gathered from open-ended questionnaires, combined with screening tools to assess the risk of mental health issues, most of these resources are focused on the English language.
+Currently, there are no open-source psychological corpora available in Romanian, and this was the main driving force behind our study, along with our interest in applying emotional analysis to such a corpus.
 3 Data Collection
-By using this survey collection methodology, participants provide their informed consent for the collection of data, as opposed to social media collected data, where data is collected without asking the participants for their consent [4].
+As opposed to social media collected data, which is collected without participants’ consent, we collected data from participants who provided their informed consent for data collection [5].
 The data collection methodology consists of a completely anonymous survey, organized in three parts.
-The first part of the survey consists of six open-ended questions.
-Its scope is to accumulate as much written text as possible.
-Out of these six questions, three of them touch positive topics and three of them negative topics.
+The first part of the survey consists of six open-ended questions, designed to gather extensive written responses.
+Out of these six questions, three touch on positive topics and three on negative topics.
 The survey was written in Romanian for Romanian speakers only.
-Below are English translations for the six open-ended questions:
+We further present the six open-ended questions, in both English and Romanian:
   1. 1.
     To what extent do you think social media affects your emotional state, especially when you see negative news?
-"În ce măsură crezi că rețelele sociale îți afectează starea emoțională, mai ales când vezi știri negative?" (negative)
+“În ce măsură crezi că rețelele sociale îți afectează starea emoțională, mai ales când vezi știri negative?” (negative)
   2. 2.
-    What is an important lesson about resilience and adaptation that you have learned recently? "Care este o lecție importantă despre reziliență și adaptare pe care ai învățat-o recent?" (positive)
+    What is an important lesson about resilience and adaptation that you have learned recently? “Care este o lecție importantă despre reziliență și adaptare pe care ai învățat-o recent?” (positive)
   3. 3.
@@
 Can you describe a time when you felt stressed because of political decisions?
-"Cât de mult crezi că politica afectează sănătatea mentală a oamenilor?
-Poți să detaliezi un moment în care ai simțit stres din cauza deciziilor politice?" (negative)
+“Cât de mult crezi că politica afectează sănătatea mentală a oamenilor?
+Poți să detaliezi un moment în care ai simțit stres din cauza deciziilor politice?” (negative)
   4. 4.
     What role do you think community or social support plays in maintaining mental health during difficult times?
-"Ce rol crezi că are comunitatea sau sprijinul social în menținerea sănătății mentale în vremuri dificile?
-"(positive)
+“Ce rol crezi că are comunitatea sau sprijinul social în menținerea sănătății mentale în vremuri dificile?” (positive)
   5. 5.
     How were you affected by discussions about war, the economic crisis, elections, or other tense topics?
-"Cum ai fost afectat de discuțiile despre război, criza economică, alegeri sau alte subiecte tensionate?" (negative)
+“Cum ai fost afectat de discuțiile despre război, criza economică, alegeri sau alte subiecte tensionate?” (negative)
   6. 6.
-    How do you recharge your batteries after a stressful period? "Cum îți reîncarci bateriile după o perioadă stresantă?" (positive)
-We wanted to create a corpus of text data written by persons prone to suffering from depression and/or anxiety.
-Usually psychological corpuses for NLP are usually gathered from social media through searching searching different keywords in social media posts that could indicate depression and/or anxiety.
-By discussing with medical professionals in the psychiatric field we concluded that this type of gathering psychological data is speculative.
-Thus, we added to our survey two more parts, one that assesses if the respondent is prone to depression by incorporating PHQ-9 [21] questions, and another part which assesses if the respondent is prone to anxiety by incorporating GAD-7 [32] questions.
-Both GAD-7 and PHQ-9 contain questions (7 for GAD-7 and 9 for PHQ-9) which ask how often the respondent has been bothered by certain problems, and the response options correspond to a numerical score: Not at all = 0, Several days = 1, More than half the days = 2, Nearly every day = 3.
+    How do you recharge your batteries after a stressful period? “Cum îți reîncarci bateriile după o perioadă stresantă?” (positive)
+Usually, psychological corpora for NLP are gathered from social media through searching for different keywords in online posts that could indicate depression and anxiety.
+However, after consulting with medical professionals in the psychiatric field, we have concluded that this method of collecting psychological data may not be entirely accurate.
+Thus, we added two more parts to our survey: one assessing respondents’ risk of depression using the PHQ-9 [23], and another assessing their risk of anxiety using the GAD-7 [35].
+Both GAD-7 and PHQ-9 contain questions (7 for GAD-7 and 9 for PHQ-9) that ask the participants how often they have experienced specific problems, and the response options correspond to a numerical score: Not at all = 0, Several days = 1, More than half the days = 2, Nearly every day = 3.
 The severity cut-offs for depression based on PHQ-9 are: 0-4: Minimal, 5-9: Mild, 10-14: Moderate, 15-19: Moderately severe, 20-27: Severe, while the severity cut-offs for anxiety based on GAD-7 are: 0-4: Minimal, 5-9: Mild, 10-14: Moderate, 15-19: Moderately severe, 20-27: Severe.
-The resulting questionnaire is completely anonymous, no personal data is requested or stored (such as name, surname, email address, IP).
-We did not collect demographic data for the same reason: to gain the trust of respondents and also to keep the total time of form completion short.
+The resulting questionnaire is completely anonymous, and no personal data is requested or stored (such as name, surname, email address, or IP).
+We did not collect demographic data for the same reason: to build trust with respondents and keep the time required to complete the form short.
 The resulting data consists of 205 responses, collected over an eight-month period from March 2025 to October 2025.
-This time frame includes two major stressful events for Romanians, presidential elections and the tax increase burden generated by the country’s new leadership.
+This time frame includes two major stressful events for Romanians - presidential elections and the tax increase burden generated by the country’s new leadership.
 4 Methodology
@@
 Based on these rules, three population groups were derived from our dataset:
-  • Depression Risk Group: 37 respondents (18%) scored higher than 10 on PHQ-9;
-  • Anxiety Risk Group: 36 respondents (17.56%) scored higher than 10 on GAD-7;
-  • No Risk Group: 28 respondents (13.66%) scored 0 for at least one out of the two questionnaires.
+  • Depression Risk Group: 37 respondents scored higher than 10 on PHQ-9 (18%);
+  • Anxiety Risk Group: 36 respondents scored higher than 10 on GAD-7 (17.56%);
+  • No Risk Group: 28 respondents scored 0 for at least one out of the two questionnaires (13.66%).
 Linguistic Inquiry and Word Count (LIWC) [6] is a lexicon and a text analysis software that provides psychologically meaningful categories and is widely used in computational mental health research [31, 24, 17, 36].
@@
 To determine which LIWC features have the greatest influence on predictions, we used SHapley Additive exPlanations (SHAP) [26].
-In parallel to these anlyses we have created an emotion detection model by fine-tuning the Romanian BERT [12] on the REDv2 dataset [8], for multi-label text classification, using the Huggingface transformers pipeline [^1] https://huggingface.co/docs/transformers.
+To complement these analyses, we have created an emotion detection model by fine-tuning the Romanian BERT [13] on the REDv2 dataset [9], for multi-label text classification, using the Huggingface transformers pipeline [^1] https://huggingface.co/docs/transformers.
 We used this model to predict the emotions in PsihoRo, thus gathering emotional information for the text written by each population group mentioned above.
@@
 Further, the cleaned DFM was transformed into the format necessary for the Structural Topic Model (STM) using which we trained the model for K = 5 topics, with the Spectral initialization method.
-In order to gain interpretable results, the words resulted from the most important topics were further expanded into their main reformulated ideas using GPT-5.
 5 Experiments and Results
@@
 max 23 21
-By analyzing the length of the answers, we questioned whether people with risk of depression and/or anxiety tend to write less.
-The answer to this question is no, because the computed T-statistic was 1.55 between the depression risk population and the no risk population, and 2.26 between the anxiety risk population and the no risk population, meaning that no significant difference was found.
+We analyzed the length of responses to determine whether individuals at risk of depression or anxiety tend to write less.
+Our findings indicate that this is not the case for Romanian.
+The computed T-statistic was 1.55 when comparing the depression risk group with the group not at risk, and 2.26 when comparing the anxiety risk group with the group not at risk.
+These results suggest that there is no significant difference in response length between these populations.
 In Figure 1 we show the distribution plots of both PHQ-9 and GAD-7 scores.
-The data showed that the distributions are skewed to the left, with the majority of observations concentrated at lower values and a longer tail extending toward the higher end of the scale, meaning that most respondents have minimal and mild risk of depression and anxiety.
+The data indicate that the distributions are skewed to the left, with the majority of observations concentrated at lower values and a longer tail extending toward the higher end of the scale, meaning that most respondents have minimal and mild risk of depression and anxiety.
 Figure 1: Distributions for GAD-7 and PHQ-9
@@
 Participants with higher PHQ-9 and GAD-7 scores discuss Achievement less frequently (-.163* and- .192**, respectively), which could indicate low motivation.
-In addition, the negative correlation with Leisure-related words (-.145* and -.169*, respectively) may stem from anhedonia, a common symptom of depression.
+In addition, the negative correlation with Leisure-related words (-.145* and -.169*, respectively) may stem from anhedonia, a common symptom of depression ([15]).
 Words in the Death category show a significant positive correlation with depression (.138*), but not with anxiety (0.029), indicating a greater preoccupation with mortality in participants with higher depression scores.
@@
 Figure 5: Emotion predictions across: all PsihoRo texts (left), pessimistic questions (middle), and optimistic questions (right).
+Figure 6: Emotion predictions across the three studied population groups.
 Surprise is an emotion that is completely absent from the PsihoRo data.
-The following emotions are more common, in sequence of appearance: anger, joy, fear, sadness, trust, and the most common, neutral.
-The emotional pattern for the more pessimistic questions is displayed in the center, while the emotional pattern for the more optimistic questions is displayed on the right.
-Emotions such as sadness and anger are on the rise on the left, while joy and trust are more prevalent on the right.
-The emotions present in the texts written by the three primary demographic groups mentioned at the start of Section 4 were also examined in Figure 6.
-It can be seen that besides the neutral, which is strongly present across all groups
-Figure 6: Emotion predictions across the three studied population groups.
-We also tried to create several regression models, based on the textual PsihoRo data gathered and the GAD-7 and PHQ-9 scores, using the Romanian BERT for text vectorization and also adding emotional information as features, but all our experiments had negative R2.
+The following emotions are more common, in order of appearance: anger, joy, fear, sadness, trust, and the most common, neutral.
+The emotional pattern for the more pessimistic questions is displayed in the center of Figure 5, while the emotional pattern for the more optimistic questions is displayed on the right.
+Emotions such as sadness and anger are on the rise on the Pessimistic Questions, while joy and trust are more prevalent on the Optimistic Questions.
+The emotions present in the texts written by the three primary demographic groups mentioned at the beginning of Section 4 are presented in Figure 6.
+While the neutral emotion is strongly represented across all groups, the population at risk for depression shows a high prevalence of sadness.
+In addition, the group at risk for anxiety exhibits emotions of sadness, fear and trust.
 5.4 Topic Modeling
-We applied the topic modeling analysis described in Section 4 for each of the three control groups studied in this article.
-The main topics are derived from the six questions asked the form.
-However, some differences appear between the three groups of respondents, which can be found in the following subsections [^2] The following subsections were generated using GPT-5, based on the words generated by the topic modeling operation, some of which present in text in bold..
+We applied the topic modeling analysis described in Section 4 for each of the three population groups studied in this work.
+The main topics are derived from the six open questions asked in the form.
+However, some differences appear among the three groups of respondents, as discussed in the following subsections.
 5.4.1 No Risk Group
-Topic 1.
-This theme highlights how individuals perceive the current political and social period in a critical and sometimes negative way.
+Topic 1 highlights how individuals perceive the current political and social period in a critical and sometimes negative way.
 The words “negative”, “decisions”, “politics”, “measure”, “news” and “period” indicate a concern for the effects of government decisions and how they are reflected in the media.
 The presence of the term “sport” suggests a possible comparison between different areas of public life.
-Topic 2.
-The theme reflects the complex interaction between information conveyed through news, political life and how they influence the general state of people.
+Topic 2 reflects the complex interaction between information conveyed through news, political life, and how they influence the general state of the population.
 Terms such as “politics”, “news”, “state”, “people” and “politics” indicate an increased attention to how media discourse shapes collective perceptions and emotions.
-Topic 3.
-This topic addresses how individuals interpret and process information related to elections and political decisions.
+Topic 3 addresses how individuals interpret and process information related to elections and political decisions.
 The terms “measure”, “depends”, “people”, “time”, “elections” and “example” suggest a personal analysis of the electoral context and an attempt to understand the relationship between information, time and decision.
-Topic 4.
-The theme highlights the importance of authentic human interaction and the balance between the inner and social dimensions of the individual.
-The words "listen", "outer", "inner", "community" and "matter" suggest an openness to empathetic communication and a valorization of social belonging.
-Topic 5.
-This topic focuses on the emotional dimension of everyday experiences, indicating how individuals manage their affective states and reactions in common situations.
-Words such as “affect”, “try”, “day”, “go” and “enough” express a constant effort to adapt and maintain psychological balance in the face of daily pressures by self regulation and personal reflection.
+Topic 4 highlights the importance of authentic human interaction and the balance between the inner and social dimensions of the individual.
+The words “listen”, “outer”, “inner”, “community” and “matter” suggest an openness to empathetic communication and highlight the value of social belonging.
+Topic 5 focuses on the emotional dimension of everyday experiences, indicating how individuals manage their affective states and reactions in common situations.
+Words such as “affect”, “try”, “day”, “go” and “enough” express a constant effort to adapt and maintain psychological balance in the face of daily pressures by self-regulation and personal reflection.
 5.4.2 Depression Risk Group
-Topic 1.
-This theme captures the perception of fatigue, daily routine and the constant effort that individuals put into their personal and professional lives.
+Topic 1 captures the perception of fatigue, daily routine and the constant effort that individuals put into their personal and professional lives.
 Associated terms such as “sleep”, “hard”, “walking” and “people” indicate a general state of exhaustion and the need for balance between personal and social life.
-Topic 2.
-This theme highlights the importance of psychological resilience and the ability to adapt to stressful or unpredictable situations.
+Topic 2 highlights the importance of psychological resilience and the ability to adapt to stressful or unpredictable situations.
 Words such as “resilience”, “situation”, “time”, “community” and “important” suggest a solution-oriented approach and maintaining a state of balance in contexts of uncertainty.
-Topic 3.
-The topic focuses on how individuals relate to political processes and the electoral climate.
+Topic 3 focuses on how individuals relate to political processes and the electoral climate.
 The terms “politics”, “elections”, “people”, “moment” and “listening” indicate both interest in political events and their influence on social attitudes and conversations.
-Topic 4.
-This topic captures the potentially harmful impact of social networks on interpersonal relationships and community cohesion.
+Topic 4 captures the potentially harmful impact of social networks on interpersonal relationships and community cohesion.
 The presence of the words “negative”, “social”, “networks”, “community” and “socialization” underlines the perception of a degradation of authentic communication and well-being caused by the digital environment.
-Topic 5.
-This theme reflects the concern for affective states and how social and economic factors contribute to the intensification of collective anxiety.
+Topic 5 reflects the concern for affective states and how social and economic factors contribute to the intensification of collective anxiety.
 The terms “feel”, “affect”, “cause”, “anxiety”, “country” and “family” suggest an emotional perspective on everyday experiences and the general social context.
 5.4.3 Anxiety Risk Group
-Topic 1.
-This topic captures how major events, such as war, economic crises or socio-political changes, affect the perception of daily life.
+Topic 1 captures how major events, such as war, economic crises or socio-political changes, affect the perception of daily life.
 Words such as “affect”, “crisis”, “war”, “elections”, “life” and “time” indicate the concern for external instability and its influence on personal balance.
-Topic 2.
-This topic highlights the close connection between political factors and individual emotional state.
+Topic 2 highlights the close connection between political factors and individual emotional state.
 Terms such as “politics”, “stress”, “affect”, “feel”, “cause” and “family” suggest a direct relationship between socio-political events and the level of anxiety or psychological tension felt.
-Topic 3.
-This topic addresses the role of community and collective resilience in the face of the constant flow of negative information from the media.
+Topic 3 addresses the role of community and collective resilience in the face of the constant flow of negative information from the media.
 The words “resilience”, “news”, “negative”, “community”, “important” and “trying” indicate an attempt to find the balance between information and protecting one’s mental state.
-Topic 4.
-This topic explores the effects of social media on psychological state, with a focus on anxiety and social pressure.
+Topic 4 explores the effects of social media on psychological state, with a focus on anxiety and social pressure.
 Terms such as “social”, “anxiety”, “state”, “hard”, “networks” and “present” suggest a connection between the intensive use of social networks and the degradation of emotional state.
-The topic reflects how virtual interactions can accentuate psychological discomfort and feelings of isolation.
-Topic 5.
-This theme emphasizes the connection between people, community and the sense of belonging.
+The topic reflects how virtual interactions can accentuate psychological discomfort.
+Topic 5 emphasizes the connection between people, community and the sense of belonging.
 The words “people”, “community”, “listen”, “dear” and “happen” highlight the importance of close relationships and solidarity in maintaining emotional balance.
 6 Conclusion and Future Work
-PsihoRo is the first open-source dataset for mental health in Romanian.
-It was created by utilizing a compound form containing both open ended questions for text gathering and the PHQ-9 and GAD-7 consecrated questionnaires to asses the risk of depression and/or anxiety of the respondents.
-Even if the dataset has only 205 entries, some very interesting insights can be observed.
-Regarding the psychologically meaningful categories from LIWC, we observed differences in categories related to body-related words, which may indicate somatic awareness, and tentative words, which indicate uncertainty, that correlated positively with both depression and anxiety scores.
+In this paper, we presented PsihoRo, the first open-source dataset for mental health in Romanian.
+The dataset was collected from 205 participants who filled in a comprehensive form that included both open-ended questions for text collection and validated questionnaires, specifically the PHQ-9 and GAD-7, to evaluate their risk of depression and anxiety.
+We analyzed the collected data using the psychologically meaningful categories from LIWC, emotion analysis and topic modeling.
+We observed differences in LIWC categories related to body-related words, which may indicate somatic awareness, and tentative words, which indicate uncertainty, that correlated positively with both depression and anxiety scores.
 A higher preoccupation with mortality was also observed in relation to depression, while anxiety scores were positively correlated with a preoccupation with family.
-While LIWC analyses showed that lexical indicators can effectively capture underlying psychological patterns related to depression and anxiety, we did not find any association between the linguistic marker ”I” and depression or anxiety in Romanian language, consistent with previous research.
-The emotion analysis showed the prevalence of sadness emotions in the population group with risk of depression, as well as fear for the risk of anxiety population group, which is also in concordance with previous research.
-The inability to develop a regression model with consistent results based on PsihoRo can be attributed both to the low number of samples and also to the naivety of assuming that depression and anxiety can be reliably discerned from texts, even when supplemented by self-assessment ratings like PHQ-9 and GAD-7.
-Depression and anxiety require meticulous analysis from multiple perspectives, rather than solely relying on an individual’s textual expression.
-However, by offering high-quality data with labels for anxiety and depression, PsihoRo seeks to support this intricate analysis.
+While LIWC analyses showed that lexical indicators can effectively capture underlying psychological patterns related to depression and anxiety, we did not find any association between the linguistic marker “I” and depression or anxiety in the Romanian language, consistent with previous research [36].
+The emotion analysis revealed a prevalence of sadness in the population group at risk for depression, as well as fear in the population group at risk for anxiety, which aligns with previous research [25, 3, 38].
+Topic modeling showed that although some perceptions and emotions are captured in textual responses, a multi-perspective approach that goes beyond self-expressed text is necessary to fully comprehend these patterns.
+By offering high-quality data with precise labels for anxiety and depression risk, PsihoRo seeks to supplement this intricate research.
 7 Ethical Considerations and Limitations
-The data collection was approved by The Ethics Committee of the University of Bucharest.
-All respondents provided their informed consent before answering the survey, which was completely anonymous, no personal data was requested or stored (such as name, surname, email address, IP, etc.).
-Adding more data by increasing the number of respondents, by using crowdfunding or similar platforms, could greatly improve PsihoRo and we aim to do this in future versions of this resource.
+The data collection received approval from the Ethics Committee of the University of Bucharest.
+All respondents provided their informed consent before participating in the survey, which was completely anonymous.
+No personal information, such as names, surnames, email addresses, or IP addresses, was requested or stored.
+We believe that increasing the number of respondents through crowdsourcing or similar platforms could significantly enhance PsihoRo. We aim to implement this in future versions of this resource.
 Acknowledgments
 The authors would like to thank psychologist Virginia Neacșu for her valuable feedback and suggestions throughout the development of this work.
-This research was supported by the University of Bucharest and Druid AI.
+This work was partially supported by Druid AI, University of Bucharest and the Ministry of Education and Research, CNCS-UEFISCDI, project SIROLA, number PN-IV-P1- PCE-2023-1701, within PNCDI IV, and by the project "Romanian Hub for Artificial Intelligence - HRIA”, Smart Growth, Digitization, and Financial Instruments Program, 2021-2027, MySMIS no.
+334906.

### v4

@@
-Alexandra Ciobotaru
+PsihoRo: Depression and Anxiety Romanian Text Corpus
 Abstract
 Psychological corpora in NLP are collections of texts used to analyze human psychology, emotions, and mental health.
-These texts allow researchers to study psychological constructs, detect mental health issues and analyze emotional language.
-However, mental health data can be difficult to collect correctly from social media, due to suppositions made by the collectors.
-A more pragmatic strategy involves gathering data through open-ended questions and then assessing this information with self-report screening surveys.
-This method was employed successfully for English, a language with a lot of psychological NLP resources.
-However, this cannot be stated for Romanian, which currently has no open-source mental health corpus.
-To address this gap, we have created the first open-source corpus for depression and anxiety in Romanian, by utilizing a form with 6 open-ended questions along with the standardized PHQ-9 and GAD-7 screening questionnaires.
-Consisting of the texts of 205 respondents and although it may seem small, PsihoRo is a first step towards understanding and analyzing texts regarding the mental health of the Romanian population.
-We employ statistical analysis, text analysis using Romanian LIWC, emotion detection and topic modeling to show what are the most important features of this newly introduced resource to the NLP community.
+These texts allow researchers to study psychological constructs, identify patterns related to mental health problems and analyze emotional language.
+However, collecting accurate mental health data from social media can be challenging due to the assumptions made by data collectors.
+A more effective approach involves gathering data through open-ended questions and then assessing participants’ mental health status using self-report screening surveys.
+This method was successfully employed for English, a language with a lot of psychological NLP resources.
+However, the same cannot be stated for Romanian, which currently has no open-source mental health corpus.
+To address this gap, we have collected the first open-source corpus focused on depression and anxiety in Romanian, by utilizing a form with 6 open-ended questions along with the standardized PHQ-9 and GAD-7 screening questionnaires.
+Although the PsihoRo corpus contains texts from only 205 respondents, it represents an important first step toward understanding and analyzing mental health issues within the Romanian population.
+We employ statistical analysis, text analysis using Romanian LIWC, emotion detection, and topic modeling to identify the most important features of this newly introduced resource for the NLP community.
 The data is publicly available at https://huggingface.co/datasets/Alegzandra/PsihoRo.
-Keywords mental health $\cdot$ Romanian $\cdot$ emotion detection $\cdot$ LIWC $\cdot$ topic modeling
-1 Introduction
+Keywords: mental health, Romanian, emotion detection, LIWC, topic modeling
+[…]
+PsihoRo: Depression and Anxiety Romanian Text Corpus
+Alexandra Ciobotaru1,2, Ana-Maria Bucur3, Liviu P. Dinu1,4
+1University of Bucharest, 2Druid AI, 3Università della Svizzera italiana,
+4Human Language Technologies Research Center
+1,490 Panduri Street, Bucharest, Romania
+21-5 Costache Negri Street, Bucharest, Romania,
+3Via Buffi 13, Lugano, Switzerland
+alexandra.zbaganu@s.unibuc.ro, ana-maria.bucur-cosma@usi.ch, ldinu@fmi.unibuc.ro
+Abstract content
+1.  Introduction
 Mental health research has increasingly used data from social media platforms, which often present a curated view of individuals’ personal lives.
-While social media can provide valuable insights for mental health studies [24, 30], there is a risk of positivity bias, where users tend to share mostly positive aspects of their experiences [32].
-Another important source for mental health research is clinical data, which typically contains sensitive information and is not publicly accessible.
+While social media can provide valuable insights for mental health studies Liu et al.
+(2022); Rai et al.
+(2024); Bucur et al.
+(2025), there is a risk of positivity bias, where users tend to share mostly positive aspects of their experiences Salopek and Eastin (2024).
+Another important source for mental health research is clinical data, which typically contains sensitive information and is often not publicly accessible Puspo et al.
+(2026).
 Most research on mental health is conducted using data in English, which may overlook important nuances present in other languages.
-Previous studies suggest that features predictive of mental health issues in English may not translate effectively to other languages or cultures [8, 30].
-For example, the use of the first-person pronoun "I" has been identified as a strong predictor of depression [31], and this feature is frequently analyzed in English-based research.
-However, the relationship between pronoun usage and depression severity can vary significantly across different demographic groups and languages [30].
-While the pronoun "I" is often used as an indicator of depression in English, its usage requires special consideration in other languages due to linguistic differences.
-For instance, previous research has shown that this association has not been observed among Romanian speakers [36].
-Romanian is a pro-drop language, which allows for ambiguity in subject pronoun usage and may contribute to a lower frequency of the personal pronoun ”I” in everyday communication.
-Despite the availability of some data in the Romanian language related to emotions, such as the RED dataset [10], REDv2 [9], and the BRIGHTER multilingual dataset [27], which includes Romanian data, there is a scarcity of mental health-related resources.
-Very few clinical data sources [36] and posts from medical forums [7, 12] exist, and those that do are not publicly accessible.
+Previous studies suggest that features predictive of mental health issues in English may not translate effectively to other languages or cultures Rai et al.
+(2024); Bucur et al.
+(2026).
+For example, the use of the first-person pronoun “I” has been identified as a strong predictor of depression Rude et al.
+(2004), and this feature is frequently analyzed in English-based research.
+However, the relationship between pronoun usage and depression severity can vary significantly across different demographic groups and languages Rai et al.
+(2024).
+While the pronoun “I” is often used as an indicator of depression in English, its use in other languages requires special consideration due to linguistic differences.
+For instance, previous research has shown that this association has not been observed among Romanian speakers Trifu et al.
+(2024).
+Romanian is a pro-drop language, which allows for ambiguity in subject pronoun usage and may contribute to a lower frequency of the personal pronoun “I” in everyday communication.
+Despite the availability of data in the Romanian language related to emotions, such as the RED dataset Ciobotaru and Dinu (2021), REDv2 Ciobotaru et al.
+(2022), and the BRIGHTER multilingual dataset Muhammad et al.
+(2025), which includes Romanian data, there is a scarcity of mental health-related resources.
+Very few clinical data sources Trifu et al.
+(2024) and posts from medical forums Briciu and Lupea (2018); Dudău and Sava (2022) exist, but unfortunately are not publicly accessible.
 This study aims to address this gap by creating the first publicly available dataset in Romanian, specifically designed to analyze emotional expressions in the context of depression and anxiety.
-By analyzing the linguistic characteristics and emotional patterns in Romanian, our research will contribute to a deeper understanding of mental health expressions in languages other than English.
+By analyzing the linguistic characteristics and emotional patterns in Romanian, our research aims to contribute to a deeper understanding of mental health expressions in languages other than English.
 Through this initiative, we seek to promote more inclusive research in mental health, paving the way for targeted interventions and support across diverse linguistic communities.
-2 Related Work
+2.  Related Work
 Previous research has gathered open-ended responses from participants to explore how expressions of mental health influence language.
-[39] created open-ended personality questions to explore how natural language processing can aid in predicting neuroticism from people’s responses.
-They gathered over 30,000 sentences written by 425 adults in Korean and used KoBERT to see which types of questions best revealed neuroticism and related traits like depression and dependency.
-[37] created the multi-modal Real World Worry Waves Dataset (RW3D), in English, which tracks the psychological effects of the COVID-19 pandemic in the UK over three consecutive years, 2020, 2021 and 2022, through repeated surveys containing also free-text responses.
-[11] used open-ended interview responses from more than 58,000 job candidates along with self-assessed personality (HEXACO) and employed “InterviewBERT” to deduce personality traits from textual replies.
-This study was also done in English.
-[33] used open-ended questionnaires to study young and middle aged adults’ responses regarding their own mental health, by employing NLP computational methods.
-They analyzed their 1106 responses and found various differences in semantics on the two study groups.
-In terms of combining emotion detection and LIWC term analysis in mental health texts, several studies employ such analysis in English: [29], [40], [2].
-[4] combined LIWC features [6], sentiment analysis and NLP to explore the relationship between language and mental health issues such as depression, anxiety, and stress during the COVID-19 pandemic.
+Yoon et al.
+(2024) created open-ended personality questions to explore how natural language processing can aid in predicting neuroticism from people’s responses.
+The authors gathered over 30,000 sentences written by 425 adults in Korean and used KoBERT to see which types of questions best revealed neuroticism and related traits like depression and dependency.
+Van der Vegt and Kleinberg (2023) created the multi-modal Real World Worry Waves Dataset (RW3D), in English, which tracks the psychological effects of the COVID-19 pandemic in the UK over three consecutive years, 2020, 2021 and 2022, through repeated surveys containing also free-text responses.
+Dai et al.
+(2022) used open-ended interview responses from more than 58,000 job candidates along with self-assessed personality (HEXACO) and employed “InterviewBERT” to deduce personality traits from textual in English.
+Sikström et al.
+(2023) used open-ended questionnaires to study young and middle-aged adults’ responses regarding their own mental health, by employing NLP computational methods.
+The authors analyzed 1,106 responses and found semantic differences between the two study groups.
+In terms of combining emotion detection and LIWC term analysis in mental health texts, several studies employ such analysis in English: Rahman et al.
+(2024), Zhang et al.
+(2023), Alshouha et al.
+(2024).
+Beech et al.
+(2025) combined LIWC features Boyd et al.
+(2022), sentiment analysis and NLP to explore the relationship between language and mental health issues such as depression, anxiety, and stress during the COVID-19 pandemic.
 They analyzed written narratives from 883 participants, alongside their DASS-21 scores, and found an association between increased symptom severity and greater use of first-person pronouns and negative emotion words.
-[21] collected responses in English from 411 participants who answered word-response questions related to their mental health and completed the GAD-7 and PHQ-9 rating scales.
+Kjell et al.
+(2021) collected responses in English from 411 participants who answered word-response questions related to their mental health and completed the GAD-7 and PHQ-9 rating scales.
 Their analyses show that individuals’ language reflects their mental health status, and combining verbal responses with quantitative measures can enhance the comprehension and assessment of the risk of depressive and anxiety symptoms in both clinical practice and research contexts.
@@
 Currently, there are no open-source psychological corpora available in Romanian, and this was the main driving force behind our study, along with our interest in applying emotional analysis to such a corpus.
-3 Data Collection
-As opposed to social media collected data, which is collected without participants’ consent, we collected data from participants who provided their informed consent for data collection [5].
+3.  Data Collection
+As opposed to social media collected data, which is collected without participants’ consent, we collected data from participants who provided their informed consent for data collection Benton et al.
+(2017).
 The data collection methodology consists of a completely anonymous survey, organized in three parts.
@@
 However, after consulting with medical professionals in the psychiatric field, we have concluded that this method of collecting psychological data may not be entirely accurate.
-Thus, we added two more parts to our survey: one assessing respondents’ risk of depression using the PHQ-9 [23], and another assessing their risk of anxiety using the GAD-7 [35].
-Both GAD-7 and PHQ-9 contain questions (7 for GAD-7 and 9 for PHQ-9) that ask the participants how often they have experienced specific problems, and the response options correspond to a numerical score: Not at all = 0, Several days = 1, More than half the days = 2, Nearly every day = 3.
+Thus, we added two more parts to our survey: one assessing respondents’ risk of depression using the PHQ-9 Kroenke et al.
+(2001), and another assessing their risk of anxiety using the GAD-7 Spitzer et al.
+(2006).
+Both GAD-7 and PHQ-9 contain questions (7 for GAD-7 and 9 for PHQ-9) that ask the participants how often they have experienced specific problems, and each item is answered on a 4-point Likert scale: Not at all = 0, Several days = 1, More than half the days = 2, Nearly every day = 3.
 The severity cut-offs for depression based on PHQ-9 are: 0-4: Minimal, 5-9: Mild, 10-14: Moderate, 15-19: Moderately severe, 20-27: Severe, while the severity cut-offs for anxiety based on GAD-7 are: 0-4: Minimal, 5-9: Mild, 10-14: Moderate, 15-19: Moderately severe, 20-27: Severe.
@@
 The resulting data consists of 205 responses, collected over an eight-month period from March 2025 to October 2025.
-This time frame includes two major stressful events for Romanians - presidential elections and the tax increase burden generated by the country’s new leadership.
-4 Methodology
-Our research consists of gathering insights from PsihoRo descriptive statistics in correlation to depression and anxiety respondents’ tendencies, for three main population groups which were computed based on the severity cut-offs.
-According to the literature in psychology [22], a cut-off score of 10 or higher on the PHQ-9 questionnaire is used to identify the presence of potential major depressive disorder, prompting further assessment.
-For the GAD-7 questionnaire, also a cut-off score of 10 has been proposed as a threshold for identifying potential generalized anxiety disorder, warranting further evaluation.
-[34].
-Based on these rules, three population groups were derived from our dataset:
+This time frame includes two major stressful events for Romanians–presidential elections and the tax increase burden generated by the country’s new leadership.
+4.  Methodology
+Our research consists of gathering insights from PsihoRo using descriptive statistics to analyze the correlation between depression and anxiety levels among respondents across three main population groups, computed based on severity cut-offs.
+According to the literature in psychology Kroenke et al.
+(2001), a PHQ-9 score of 10 or higher indicates the potential presence of major depressive disorder and prompts further assessment.
+For the GAD-7 questionnaire, a cut-off score of 10 has also been proposed as a threshold for identifying potential generalized anxiety disorder, warranting further evaluation Spitzer et al.
+(2006).
+Based on these cut-off scores, three population groups were derived from our dataset:
   • Depression Risk Group: 37 respondents scored higher than 10 on PHQ-9 (18%);
@@
   • No Risk Group: 28 respondents scored 0 for at least one out of the two questionnaires (13.66%).
-Linguistic Inquiry and Word Count (LIWC) [6] is a lexicon and a text analysis software that provides psychologically meaningful categories and is widely used in computational mental health research [31, 24, 17, 36].
-The official LIWC software is available for analyzing textual data in the English language; however, validated translations of the software are also available in several languages.
-In our experiments, we use the Romanian version of LIWC (Ro-LIWC2015) [12].
+Linguistic Inquiry and Word Count (LIWC) Boyd et al.
+(2022) is a lexicon and a text analysis software that provides psychologically meaningful categories and is widely used in computational mental health research Rude et al.
+(2004); Liu et al.
+(2022); Hur et al.
+(2024); Trifu et al.
+(2024).
+The official LIWC software is available for analyzing textual data in the English language; however, validated translations of the lexicon are also available in several languages.
+In our experiments, we use the Romanian version of LIWC (Ro-LIWC2015) Dudău and Sava (2022).
 The 2015 version of LIWC includes 98 categories that relate to linguistic dimensions and grammar, as well as psychological processes, including affective, cognitive, social, perceptual, and biological processes.
-Previous research has indicated that individuals with mental health issues tend to exhibit specific language patterns.
-These patterns include a higher frequency of using the personal pronoun "I" [31], increased use of words with negative connotations [14], and a tendency to use words with absolute meanings [1].
-These findings highlight the importance of language analysis in understanding mental health and motivate us to use LIWC to examine the language patterns of the participants.
-After checking for differences in LIWC categories in relation to the PHQ-9 and GAD-7 scores, we wanted to test their discriminatory power in classification.
-To this end, we employed a LightGBM classifier [20] and trained it on the results with LIWC to predict a binary label for depression and anxiety, based on the cutoff score of 10.
-The model achieved an accuracy of 0.85 in classifying depression and 0.83 in classifying anxiety.
-To determine which LIWC features have the greatest influence on predictions, we used SHapley Additive exPlanations (SHAP) [26].
-To complement these analyses, we have created an emotion detection model by fine-tuning the Romanian BERT [13] on the REDv2 dataset [9], for multi-label text classification, using the Huggingface transformers pipeline [^1] https://huggingface.co/docs/transformers.
-We used this model to predict the emotions in PsihoRo, thus gathering emotional information for the text written by each population group mentioned above.
-The resulting model had an F1 score of 66.85% computed on the REDv2 test set.
-Another interesting analysis was topic modeling.
-To identify the main themes in the PsihoRo corpus we used the stm R package.
+Previous research on the English language has indicated that individuals with mental health issues tend to exhibit specific language patterns.
+These patterns include a higher frequency of using the personal pronoun “I” Rude et al.
+(2004), increased use of words with negative connotations Fekete (2002), and a tendency to use words with absolute meanings Al-Mosaiwi and Johnstone (2018).
+These findings highlight the importance of language analysis for understanding mental health and motivate us to use LIWC to examine participants’ language patterns.
+After checking for differences in LIWC categories across PHQ-9 and GAD-7 scores, we wanted to test their discriminatory power for classification.
+To this end, we employed a LightGBM classifier Ke et al.
+(2017) and trained it on LIWC results to predict a binary label for depression and anxiety, using a cut-off score of 10.
+The model achieved accuracies of 0.85 for depression and 0.83 for anxiety.
+To determine which LIWC features have the greatest influence on predictions, we used SHapley Additive exPlanations (SHAP) Lundberg and Lee (2017).
+To complement these analyses, we have created an emotion detection model by fine-tuning the Romanian BERT Dumitrescu et al.
+(2020) on the REDv2 dataset Ciobotaru et al.
+(2022), for multi-label text classification, using the Huggingface transformers pipeline.
+[^1] https://huggingface.co/docs/transformers We used this model to predict the emotions in PsihoRo, thus gathering emotional information for the text written by each population group mentioned above.
+The resulting model achieved an F1 score of 66.85% on the REDv2 test set.
+We also performed topic modeling to identify the main themes in the PsihoRo corpus, using the stm R package.
 First, the texts were tokenized, lowercased, and cleaned by removing punctuation, numbers, symbols and stopwords.
 Next, we created a Document-Feature Matrix (DFM) to represent term frequencies across texts.
-Also, to enhance model performance and reduce noise, we excluded rare terms which appeared in fewer than 3 documents and the very common terms, which appeared in more than 90% of the texts.
-Further, the cleaned DFM was transformed into the format necessary for the Structural Topic Model (STM) using which we trained the model for K = 5 topics, with the Spectral initialization method.
-5 Experiments and Results
-5.1 PsihoRo Descriptive Statistics
-In Table 1 we show mean, median (50% quantile), min, max, 25% and 75% cuantiles, and the standard deviation for both anxiety and depression scores.
-It can be seen that the maximum value for GAD-7 (anxiety) is 21, which is the maximum which can be achieved, and the maximum score for PHQ-9 (depression) is 23, out of the maximum possible of 27.
-Table 1: Descriptive statistics for PsihoRo
+To enhance model performance and reduce noise, we excluded rare terms that appeared in fewer than 3 documents and the very common terms that appeared in more than 90% of the texts.
+Further, the cleaned DFM was transformed into the format required by the Structural Topic Model (STM), and we trained the model with $K$ = 5 topics using the Spectral initialization method.
+5.  Experiments and Results
+5.1.  PsihoRo Descriptive Statistics
+In Table 1, we show mean, median (50% percentile), min, max, 25% and 75% percentiles, and the standard deviation for both anxiety and depression scores.
+In the collected data, the maximum value for GAD-7 (anxiety) is 21, which is the maximum that can be achieved, and the maximum score for PHQ-9 (depression) is 23, out of the maximum possible of 27.
 PHQ-9 GAD-7
 Count 205 205
-mean 6 5.78
-standard deviation 4.83 4.95
+Mean 6.00 5.78
+SD 4.83 4.95
 Min 0 0
@@
 Max 23 21
+Table 1: Descriptive statistics for PsihoRo
 We analyzed the length of responses to determine whether individuals at risk of depression or anxiety tend to write less.
@@
 These results suggest that there is no significant difference in response length between these populations.
-In Figure 1 we show the distribution plots of both PHQ-9 and GAD-7 scores.
-The data indicate that the distributions are skewed to the left, with the majority of observations concentrated at lower values and a longer tail extending toward the higher end of the scale, meaning that most respondents have minimal and mild risk of depression and anxiety.
+In Figure 1, we present the distribution plots of both PHQ-9 and GAD-7 scores.
+The data indicate that the distributions are skewed to the left, with most observations clustered at lower values and a longer tail extending toward the higher end of the scale.
+This suggests that the majority of respondents experience minimal to mild risk of depression and anxiety.
 Figure 1: Distributions for GAD-7 and PHQ-9
-The Pearson correlation coefficient (r) between PHQ-9 and GAD-7 total score is 0.761 (p<0.001), which means that anxiety and depression are strongly correlated in our data.
-This result shows as well in Figure 2 and is in concordance with the results of several other works: [19], [16], [18].
+The Pearson correlation coefficient ($r$) between the PHQ-9 and GAD-7 total scores is 0.761 (pp<0.001), indicating that anxiety and depression are strongly correlated in our data.
+This finding is illustrated in Figure 2 and aligns with the results of several other studies Kalin (2020); Hoying et al.
+(2020); Jacobson and Newman (2014).
 Figure 2: Correlation between Anxiety and Depression Scores.
-5.2 LIWC
-We employ LIWC to analyze language patterns in participants’ open-ended responses and correlate them with scores from the PHQ-9 and GAD-7.
-In Table 2, we present the Pearson correlation coefficients between LIWC categories and the symptom scores.
-Positive values indicate that higher symptom scores are associated with greater use of that linguistic category, while negative values indicate the opposite.
-Asterisks indicate significance (*p < .05, **p < .01).
-Surprisingly, the use of negations is negatively correlated with both PHQ-9 and GAD-7 scores, suggesting that participants with higher depression and anxiety scores tend to use fewer negations in their responses.
-Individuals who score higher on these assessments tend to use more words in the Tentative category, which indicates uncertainty, supported by the correlation coefficients of .204** for PHQ-9 and .213** for GAD-7.
-Comparatives also show a positive correlation with the depression score (.167*), indicating comparative thinking, which is often linked to depressive rumination.
-However, another marker for rumination, the high frequency of Past focus verbs, is not observed in our data.
-There is a positive correlation between the use of Body-related words and PHQ-9 (200**) and GAD-7 (180**) scores, indicating that participants with higher scores frequently use more language related to bodily sensations.
-This may reflect increased somatic awareness, a common trait in depression and anxiety [28].
-Participants with higher PHQ-9 and GAD-7 scores discuss Achievement less frequently (-.163* and- .192**, respectively), which could indicate low motivation.
-In addition, the negative correlation with Leisure-related words (-.145* and -.169*, respectively) may stem from anhedonia, a common symptom of depression ([15]).
-Words in the Death category show a significant positive correlation with depression (.138*), but not with anxiety (0.029), indicating a greater preoccupation with mortality in participants with higher depression scores.
-In contrast, the Family category is positively correlated with anxiety scores (r = 0.145*), suggesting that individuals with higher anxiety are more focused on family-related issues.
-Table 2: Correlations between PHQ-9, GAD-7, and LIWC dimensions
-LIWC Dimension Examples PHQ-9 GAD-7
-I I, me, mine 0.018 -0.046
-We we, us, our 0.013 -0.039
-You you, your 0.006 0.084
-Verb eat, come, carry -0.087 -0.091
-Past focus ago, did, talked -0.064 -0.075
-Present focus today, is, now -0.019 -0.034
-Future focus may, will, soon -0.027 -0.055
-Adjectives free, happy, long 0.118 0.12
-Negations no, not, never -.156* -.167*
-Positive emotion love, nice, sweet -0.122 -0.057
-Negative emotion hurt, ugly, nasty 0.028 0.028
-Anxiety worried, fearful -0.042 0.01
-Anger hate, kill, annoyed 0.107 0.105
-Sadness crying, grief, sad 0.074 0.045
-Family daughter, dad, aunt 0.079 .145*
-Friends buddy, neighbor -0.026 0.043
-Female references girl, her, mom 0.061 0.074
-Male references boy, his, dad .156* .177*
-Comparatives greater, best, after .167* 0.081
-Causation because, effect -.186** -.180**
-Tentative maybe, perhaps .204** .213**
-Insight think, know -0.075 -0.101
-Assent agree, OK, yes 0.009 0.065
-Perceptual processes look, heard, feeling 0.009 0.033
-Body cheek, hands, spit .200** .180**
-Health clinic, flu, pill -0.051 0.035
-Achievement win, success, better -.163* -.192**
-Leisure cook, chat, movie -.145* -.169*
-Death bury, coffin, kill .138* 0.029
+LIWC Dimension Examples PHQ-9 GAD-7 I I, me, mine .018 -.046 We we, us, our .013 -.039 You you, your .006 .084 Verb eat, come, carry -.087 -.091 Past focus ago, did, talked -.064 -.075 Present focus today, is, now -.019 -.034 Future focus may, will, soon -.027 -.055 Adjectives free, happy, long .118 .120 Negations no, not, never -.156* -.167* Positive emotion love, nice, sweet -.122 -.057 Negative emotion hurt, ugly, nasty .028 .028 Anxiety worried, fearful -.042 .010 Anger hate, kill, annoyed .107 .105 Sadness crying, grief, sad .074 .045 Family daughter, dad, aunt .079 .145* Friends buddy, neighbor -.026 .043 Female references girl, her, mom .061 .074 Male references boy, his, dad .156* .177* Comparatives greater, best, after .167* .081 Causation because, effect -.186** -.180** Tentative maybe, perhaps .204** .213** Insight think, know -.075 -.101 Assent agree, OK, yes .009 .065 Perceptual processes look, heard, feeling .009 .033 Body cheek, hands, spit .200** .180** Health clinic, flu, pill -.051 .035 Achievement win, success, better -.163* -.192** Leisure cook, chat, movie -.145* -.169* Death bury, coffin, kill .138* .029
 Note. * $p<.05$, ** $p<.01$.
+Table 2: Correlations between PHQ-9, GAD-7, and LIWC dimensions.
 Figure 3: Differences in selected LIWC linguistic categories between participants with low vs.
@@
 Positive values indicate greater relative use among participants with higher symptom scores.
+5.2.  LIWC
+We employ LIWC to analyze language patterns in participants’ open-ended responses and correlate them with scores from the PHQ-9 and GAD-7.
+In Table 2, we present the Pearson correlation coefficients between LIWC categories and the symptom scores.
+Positive values indicate that higher symptom scores are associated with greater use of that linguistic category, while negative values indicate the opposite.
+Figure 4: Feature importance for LIWC categories analyzed using SHAP.
+Surprisingly, the use of negations is negatively correlated with both PHQ-9 and GAD-7 scores, suggesting that participants with higher depression and anxiety scores tend to use fewer negations in their responses.
+Individuals who score higher on the PHQ-9 and GAD-7 tend to use more words in the Tentative category, which indicates uncertainty, as evidenced by correlation coefficients of .204** for PHQ-9 and .213** for GAD-7.
+Comparatives also show a positive correlation with the depression score (.167*), indicating comparative thinking, which is often linked to depressive rumination Nolen-Hoeksema et al.
+(2008).
+However, another marker for rumination, the high frequency of Past focus verbs, is not observed in our data.
+There is a positive correlation between the use of Body-related words and PHQ-9 (.200**) and GAD-7 (.180**) scores, indicating that participants with higher scores frequently use more language related to bodily sensations.
+This may reflect increased somatic awareness, a common trait in depression and anxiety Penninx et al.
+(2013).
+Participants with higher PHQ-9 and GAD-7 scores discuss Achievement less frequently (-.163* and -.192**, respectively), which could indicate lower motivation.
+In addition, the negative correlation with Leisure-related words (-.145* and -.169*, respectively) may stem from anhedonia, a common symptom of depression Grillo (2016).
+Words in the Death category show a significant positive correlation with depression (.138*), but not with anxiety, indicating a greater preoccupation with mortality in participants with higher depression scores.
+In contrast, the Family category is positively correlated with anxiety scores (r = .145*), suggesting that individuals with higher anxiety are more focused on family-related issues.
 To better understand the differences between higher and lower symptom scores for depression and anxiety, we conducted an additional analysis of the LIWC categories.
-According to the cutoff values for both the PHQ-9 and GAD-7 assessment tools, we define a low PHQ score as anything less than 10, while a score greater than 10 is classified as a high PHQ-9 score.
+According to the cut-off values for both the PHQ-9 and GAD-7, we define a low PHQ-9 score as less than 10, while a score of 10 or higher is considered high.
 In Figure 3, the goal was to identify linguistic markers that differentiate individuals with higher vs.
@@
 The analysis produced results consistent with the correlation analysis, particularly regarding Achievement and tentativeness in relation to depression and anxiety.
-However, for depression, we observed a decreased use of Positive emotion words and an increased usage of Sadness words, which aligns with existing literature in psychology [36].
+However, for depression, we observed a decreased use of Positive emotion words and an increased usage of Sadness words, which aligns with existing literature in psychology Trifu et al.
+(2024).
 In addition, the language of participants with high GAD-7 scores showed a higher frequency of words in the Interrogatives category, indicating a tendency to question or seek clarity.
-Figure 4: Feature importance for LIWC categories analyzed using SHAP.
-To determine which LIWC features have the greatest influence on predictions, we used SHapley Additive exPlanations (SHAP) [26], as explained in Section 4, with the results presented in Figure 4.
+To determine which LIWC features have the greatest influence on predictions, we used SHapley Additive exPlanations (SHAP) Lundberg and Lee (2017), as explained in Section 4, with the results presented in Figure 4.
 The importance of the top 20 features aligns with our correlation analysis and highlights the differences between low and high scores for both depression and anxiety.
+For predicting depression, low values in the LIWC features for Assent, Verb, and Common Adjectives, along with high values in Tentative words, contribute to the model’s positive predictions.
+In the case of anxiety, low values in LIWC features related to Achievement and Insight, combined with high values in Function words, Interrogatives, and Conjunctions, also have a higher contribution to a positive prediction.
 The LIWC analyses presented in this section demonstrate that lexical indicators can effectively capture underlying psychological patterns related to depression and anxiety.
-However, the most important feature used in mental health research, the pronoun ”I”, appears to be less useful for the Romanian language, likely because it is a pro-drop language.
-This finding is consistent with the work of [36], which also indicated that this marker does not generalize well in a sample of patients with major depression.
-5.3 Emotion Analysis
-Using the emotion detection model described in Section 4, we predicted the emotions in all the texts of PsihoRo, and then we aggregated the results by using radar charts to visualize the overall emotional patterns.
-In Figure 5 we show the emotion predictions across all PsihoRo texts, and also for the pessimistic questions (1, 3, 5) and the optimistic ones (2, 4, 6).
-Figure 5: Emotion predictions across: all PsihoRo texts (left), pessimistic questions (middle), and optimistic questions (right).
+However, the most important feature used in mental health research, the pronoun “I”, appears to be less useful for the Romanian language, likely because it is a pro-drop language.
+This finding is consistent with the work of Trifu et al.
+(2024), which also indicated that this marker does not generalize well in a sample of Romanian patients with major depression.
+5.3.  Emotion Analysis
+Using the emotion detection model described in Section 4, we predicted emotions for all PsihoRo texts, then aggregated the results using radar charts to visualize overall emotional patterns.
+In Figure 5, we show the emotion predictions across all PsihoRo texts, as well as for the negative-topic (pessimistic) questions (1, 3, 5) and the positive-topic (optimistic) questions (2, 4, 6).
+Figure 5: Emotion predictions across: all PsihoRo texts (left), negative-topic (pessimistic) questions (middle), and positive-topic (optimistic) questions (right).
 Figure 6: Emotion predictions across the three studied population groups.
 Surprise is an emotion that is completely absent from the PsihoRo data.
-The following emotions are more common, in order of appearance: anger, joy, fear, sadness, trust, and the most common, neutral.
+The following emotions are more common, in order of appearance: anger, joy, fear, sadness, trust, and, most commonly, neutral.
 The emotional pattern for the more pessimistic questions is displayed in the center of Figure 5, while the emotional pattern for the more optimistic questions is displayed on the right.
@@
 In addition, the group at risk for anxiety exhibits emotions of sadness, fear and trust.
-5.4 Topic Modeling
+5.4.  Topic Modeling
 We applied the topic modeling analysis described in Section 4 for each of the three population groups studied in this work.
@@
 However, some differences appear among the three groups of respondents, as discussed in the following subsections.
-5.4.1 No Risk Group
+5.4.1.  No Risk Group
 Topic 1 highlights how individuals perceive the current political and social period in a critical and sometimes negative way.
@@
 Words such as “affect”, “try”, “day”, “go” and “enough” express a constant effort to adapt and maintain psychological balance in the face of daily pressures by self-regulation and personal reflection.
-5.4.2 Depression Risk Group
+5.4.2.  Depression Risk Group
 Topic 1 captures the perception of fatigue, daily routine and the constant effort that individuals put into their personal and professional lives.
@@
 The terms “feel”, “affect”, “cause”, “anxiety”, “country” and “family” suggest an emotional perspective on everyday experiences and the general social context.
-5.4.3 Anxiety Risk Group
+5.4.3.  Anxiety Risk Group
 Topic 1 captures how major events, such as war, economic crises or socio-political changes, affect the perception of daily life.
@@
 The words “people”, “community”, “listen”, “dear” and “happen” highlight the importance of close relationships and solidarity in maintaining emotional balance.
-6 Conclusion and Future Work
+6.  Conclusion and Future Work
 In this paper, we presented PsihoRo, the first open-source dataset for mental health in Romanian.
@@
 A higher preoccupation with mortality was also observed in relation to depression, while anxiety scores were positively correlated with a preoccupation with family.
-While LIWC analyses showed that lexical indicators can effectively capture underlying psychological patterns related to depression and anxiety, we did not find any association between the linguistic marker “I” and depression or anxiety in the Romanian language, consistent with previous research [36].
-The emotion analysis revealed a prevalence of sadness in the population group at risk for depression, as well as fear in the population group at risk for anxiety, which aligns with previous research [25, 3, 38].
+While LIWC analyses showed that lexical indicators can effectively capture underlying psychological patterns related to depression and anxiety, we did not find any association between the linguistic marker “I” and depression or anxiety in the Romanian language, consistent with previous research Trifu et al.
+(2024).
+The emotion analysis revealed a prevalence of sadness in the population group at risk for depression, as well as fear in the population group at risk for anxiety, which aligns with previous research Low et al.
+(2020); Atapattu et al.
+(2022); Varastehnezhad et al.
+(2025).
 Topic modeling showed that although some perceptions and emotions are captured in textual responses, a multi-perspective approach that goes beyond self-expressed text is necessary to fully comprehend these patterns.
 By offering high-quality data with precise labels for anxiety and depression risk, PsihoRo seeks to supplement this intricate research.
-7 Ethical Considerations and Limitations
+Ethical Considerations and Limitations
 The data collection received approval from the Ethics Committee of the University of Bucharest.
@@
 334906.
+Bibliographical References
+  • M.
+Al-Mosaiwi and T.
+Johnstone (2018) In an absolute state: elevated use of absolutist words is a marker specific to anxiety, depression, and suicidal ideation.
+Clinical psychological science 6 (4), pp.
+529–542.
+Cited by: §4.
+  • B.
+Alshouha, J.
+Serrano‐Guerrero, F.
+Chiclana, F.
+P.
+Romero, and J.
+A.
+Olivas (2024) BioEmoDetector: a flexible platform for detecting emotions from health narratives.
+pp.
+101670.
+External Links: Document, Link Cited by: §2.
+  • T.
+Atapattu, M.
+Herath, C.
+Elvitigala, P.
+de Zoysa, K.
+Gunawardana, M.
+Thilakaratne, K.
+de Zoysa, and K.
+Falkner (2022) EmoMent: an emotion annotated mental health corpus from two South Asian countries.
+In Proceedings of the 29th International Conference on Computational Linguistics, N.
+Calzolari, C.
+Huang, H.
+Kim, J.
+Pustejovsky, L.
+Wanner, K.
+Choi, P.
+Ryu, H.
+Chen, L.
+Donatelli, H.
+Ji, S.
+Kurohashi, P.
+Paggio, N.
+Xue, S.
+Kim, Y.
+Hahm, Z.
+He, T.
+K.
+Lee, E.
+Santus, F.
+Bond, and S.
+Na (Eds.), Gyeongju, Republic of Korea, pp.
+6991–7001.
+External Links: Link Cited by: §6.
+  • A.
+Beech, H.
+Fan, J.
+Shu, J.
+Oyarzun, P.
+Nadel, O.
+T.
+Karaman, S.
+Vranos, E.
+A.
+Phelps, and M.
+A.
+Kredlow (2025) Using natural language processing to identify patterns associated with depression, anxiety, and stress symptoms during the covid-19 pandemic.
+376, pp.
+113–121.
+External Links: ISSN 0165-0327, Document, Link Cited by: §2.
+  • A.
+Benton, G.
+Coppersmith, and M.
+Dredze (2017) Ethical research protocols for social media health research.
+In Proceedings of the first ACL workshop on ethics in natural language processing, pp.
+94–102.
+Cited by: §3.
+  • R. L. Boyd, A. Ashokkumar, S. Seraj, and J. W. Pennebaker (2022) The development and psychometric properties of liwc-22. Austin, TX: University of Texas at Austin 10, pp. 1–47. Cited by: §2, §4.
+  • A.
+Briciu and M.
+Lupea (2018) Studying the language of mental illness in romanian social media.
+In 2018 IEEE 14th International Conference on Intelligent Computer Communication and Processing (ICCP), pp.
+21–28.
+Cited by: §1.
+  • A.
+Bucur, A.
+Moldovan, K.
+Parvatikar, M.
+Zampieri, A.
+R.
+KhudaBukhsh, and L.
+P.
+Dinu (2025) On the state of nlp approaches to modeling depression in social media: a post-covid-19 outlook.
+IEEE Journal of Biomedical and Health Informatics, pp.
+1–13.
+External Links: Document Cited by: §1.
+  • A.
+Bucur, M.
+Zampieri, T.
+Ranasinghe, and F.
+Crestani (2026) A survey on multilingual mental disorders detection from social media data.
+In Proceedings of the 19th Conference of the European Chapter of the Association for Computational Linguistics (Volume 1: Long Papers), Cited by: §1.
+  • A.
+Ciobotaru, M.
+V.
+Constantinescu, L.
+P.
+Dinu, and S.
+Dumitrescu (2022) RED v2: enhancing RED dataset for multi-label emotion detection.
+In Proceedings of the Thirteenth Language Resources and Evaluation Conference, N.
+Calzolari, F.
+Béchet, P.
+Blache, K.
+Choukri, C.
+Cieri, T.
+Declerck, S.
+Goggi, H.
+Isahara, B.
+Maegaard, J.
+Mariani, H.
+Mazo, J.
+Odijk, and S.
+Piperidis (Eds.), Marseille, France, pp.
+1392–1399.
+External Links: Link Cited by: §1, §4.
+  • A.
+Ciobotaru and L.
+P.
+Dinu (2021) RED: a novel dataset for romanian emotion detection from tweets.
+In Proceedings of the international conference on recent advances in natural language processing (RANLP 2021), pp.
+291–300.
+Cited by: §1.
+  • Y. Dai, M. Jayaratne, and B. Jayatilleke (2022) Explainable personality prediction using answers to open-ended interview questions. 13. External Links: Document, Link, ISSN 1664-1078 Cited by: §2.
+  • D.
+P.
+Dudău and F.
+A.
+Sava (2022) The development and validation of the romanian version of linguistic inquiry and word count 2015 (ro-liwc2015).
+Current Psychology 41 (6), pp.
+3597–3614.
+Cited by: §1, §4.
+  • S.
+Dumitrescu, A.
+Avram, and S.
+Pyysalo (2020) The birth of Romanian BERT.
+In Findings of the Association for Computational Linguistics: EMNLP 2020, T.
+Cohn, Y.
+He, and Y.
+Liu (Eds.), Online, pp.
+4324–4328.
+External Links: Link, Document Cited by: §4.
+  • S. Fekete (2002) The internet-a new source of data on suicide, depression and anxiety: a preliminary study. Archives of Suicide Research 6 (4), pp. 351–361. Cited by: §4.
+  • L. Grillo (2016) A possible role of anhedonia as common substrate for depression and anxiety. 2016, pp. 1598130. External Links: Document Cited by: §5.2.
+  • J.
+Hoying, B.
+M.
+Melnyk, E.
+Hutson, and A.
+Tan (2020) Prevalence and correlates of depression, anxiety, stress, healthy beliefs, and lifestyle behaviors in first-year graduate health sciences students.
+17 (1), pp.
+49–59.
+Note: Epub 2020 Jan 8 External Links: Document, Link Cited by: §5.1.
+  • J. K. Hur, J. Heffner, G. W. Feng, J. Joormann, and R. B. Rutledge (2024) Language sentiment predicts changes in depressive symptoms. 121 (39), pp. e2321321121. Cited by: §4.
+  • N.
+C.
+Jacobson and M.
+G.
+Newman (2014) Avoidance mediates the relationship between anxiety and depression over a decade later.
+28 (5), pp.
+437–445.
+Note: Epub 2014 Apr 16 External Links: Document, Link Cited by: §5.1.
+  • N.
+H.
+Kalin (2020) The critical relationship between anxiety and depression.
+177 (5), pp.
+365–367.
+Note: PMID: 32354270 External Links: Document, Link, https://doi.org/10.1176/appi.ajp.2020.20030305 Cited by: §5.1.
+  • G.
+Ke, Q.
+Meng, T.
+Finley, T.
+Wang, W.
+Chen, W.
+Ma, Q.
+Ye, and T.
+Liu (2017) Lightgbm: a highly efficient gradient boosting decision tree.
+Advances in neural information processing systems 30.
+Cited by: §4.
+  • O.
+N.
+E.
+Kjell, P.
+Johnsson, and S.
+Sikström (2021) Freely generated word responses analyzed with artificial intelligence predict self-reported symptoms of depression, anxiety, and worry.
+12, pp.
+602581.
+External Links: Document, Link Cited by: §2.
+  • K.
+Kroenke, R.
+L.
+Spitzer, and J.
+B.
+W.
+Williams (2001) The phq-9: validity of a brief depression severity measure.
+Journal of General Internal Medicine 16 (9), pp.
+606–613.
+External Links: Document Cited by: §3, §4.
+  • T.
+Liu, L.
+H.
+Ungar, B.
+Curtis, G.
+Sherman, K.
+Yadeta, L.
+Tay, J.
+C.
+Eichstaedt, and S.
+C.
+Guntuku (2022) Head versus heart: social media reveals differential language of loneliness from depression.
+1 (1), pp.
+16.
+Cited by: §1, §4.
+  • D.
+M.
+Low, L.
+Rumker, T.
+Talkar, J.
+Torous, G.
+Cecchi, and S.
+S.
+Ghosh (2020) Natural language processing reveals vulnerable mental health support groups and heightened health anxiety on reddit during covid-19: observational study.
+22 (10), pp.
+e22635.
+External Links: Document Cited by: §6.
+  • S. M. Lundberg and S. Lee (2017) A unified approach to interpreting model predictions. Advances in neural information processing systems 30. Cited by: §4, §5.2.
+  • S.
+H.
+Muhammad, N.
+Ousidhoum, I.
+Abdulmumin, J.
+P.
+Wahle, T.
+Ruas, et al.
+(2025) BRIGHTER: BRIdging the gap in human-annotated textual emotion recognition datasets for 28 languages.
+In Proceedings of the 63rd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), W.
+Che, J.
+Nabende, E.
+Shutova, and M.
+T.
+Pilehvar (Eds.), Vienna, Austria, pp.
+8895–8916.
+External Links: Link, Document, ISBN 979-8-89176-251-0 Cited by: §1.
+  • S. Nolen-Hoeksema, B. E. Wisco, and S. Lyubomirsky (2008) Rethinking rumination. Perspectives on psychological science 3 (5), pp. 400–424. Cited by: §5.2.
+  • B.
+W.
+Penninx, Y.
+Milaneschi, F.
+Lamers, and N.
+Vogelzangs (2013) Understanding the somatic consequences of depression: biological mechanisms and the role of depression symptom profile.
+BMC medicine 11 (1), pp.
+129.
+Cited by: §5.2.
+  • S.
+S.
+C.
+Puspo, A.
+Bucur, S.
+Chancellor, O.
+Uzuner, and M.
+Zampieri (2026) Mental health disorder detection beyond social media: a systematic review of available datasets.
+In Proceedings of LREC, Cited by: §1.
+  • A.
+B.
+S.
+Rahman, H.
+Ta, L.
+Najjar, A.
+Azadmanesh, and A.
+S.
+Gönül (2024) DepressionEmo: a novel dataset for multilabel classification of depression emotions.
+Note: Preprint External Links: Link Cited by: §2.
+  • S.
+Rai, E.
+C.
+Stade, S.
+Giorgi, A.
+Francisco, L.
+H.
+Ungar, B.
+Curtis, and S.
+C.
+Guntuku (2024) Key language markers of depression on social media depend on race.
+Proceedings of the National Academy of Sciences 121 (14), pp.
+e2319837121.
+Cited by: §1, §1.
+  • S. Rude, E. Gortner, and J. Pennebaker (2004) Language use of depressed and depression-vulnerable college students. Cognition & Emotion 18 (8), pp. 1121–1133. Cited by: §1, §4, §4.
+  • A.
+H.
+Salopek and M.
+S.
+Eastin (2024) Toxic positivity intentions: an image management approach to upward social comparison and false self-presentation.
+Journal of Computer-Mediated Communication 29 (3), pp.
+zmae003.
+Cited by: §1.
+  • S.
+Sikström, B.
+Kelmendi, and N.
+Persson (2023) Assessment of depression and anxiety in young and old with a question-based computational language approach.
+2, pp.
+11.
+External Links: Document, Link Cited by: §2.
+  • R.
+L.
+Spitzer, K.
+Kroenke, J.
+B.
+W.
+Williams, and B.
+Löwe (2006) A brief measure for assessing generalized anxiety disorder: the gad-7.
+Archives of Internal Medicine 166 (10), pp.
+1092–1097.
+External Links: Document Cited by: §3, §4.
+  • R.
+N.
+Trifu, B.
+Nemeș, D.
+C.
+Herta, C.
+Bodea-Hategan, D.
+A.
+Talaș, and H.
+Coman (2024) Linguistic markers for major depressive disorder: a cross-sectional study using an automated procedure.
+Frontiers in Psychology 15, pp.
+1355734.
+Cited by: §1, §1, §4, §5.2, §5.2, §6.
+  • I.
+Van der Vegt and B.
+Kleinberg (2023) A multi-modal panel dataset to understand the psychological impact of the pandemic.
+Scientific Data 10 (1), pp.
+537.
+External Links: Document, Link, ISSN 2052-4463 Cited by: §2.
+  • A.
+Varastehnezhad, R.
+Tavasoli, S.
+Elyasi, M.
+LotfiNia, and H.
+Farbeh (2025) AI in mental health: emotional and sentiment analysis of large language models’ responses to depression, anxiety, and stress queries.
+abs/2508.11285.
+External Links: Link Cited by: §6.
+  • S.
+Yoon, J.
+Jang, G.
+Son, S.
+Park, J.
+Hwang, J.
+Y.
+Choeh, and K.
+Choi (2024) Predicting neuroticism with open-ended response using natural language processing.
+Frontiers in Psychiatry 15, pp.
+1437569.
+External Links: Document, Link Cited by: §2.
+  • T.
+Zhang, K.
+Yang, S.
+Ji, and S.
+Ananiadou (2023) Emotion fusion for mental illness detection from social media: a survey.
+92, pp.
+231–246.
+External Links: ISSN 1566-2535, Document, Link Cited by: §2.